The molar Sr/Ca ratio of aragonite (CaCO3) is 8.5 millimoles/mole, what is the concentration of Sr in ppm?
1
Expert's answer
2020-04-18T07:07:57-0400

Mass of Sr is 0.0085*87.62 = 0.7448 g. Concentration of Sr = 0.7448/101.34 = 0.00735 and concentration in ppm = 0.00735 * 106 = 7350 PPM.